OpenStudy (anonymous):

Two angles are vertical angles one measures (2x+6)° and the other measures 80° find the value of x?

OpenStudy (anonymous):

You clearly do not like me trying to answer this question, however I will do so anyway. I do not understand why you were given this problem as you do not know how to substitute, so I will do it for you. After substitution, you get 2x-6=80. You can solve that at least I hope

OpenStudy (anonymous):

??

OpenStudy (anonymous):

what is 2x-6=80

OpenStudy (anonymous):

If you cannot do that then give up on solving this question, and tell your teacher or whatever else gave you this problem and say you cant do this problem

OpenStudy (anonymous):

6 + 2x = 80 Solving 6 + 2x = 80 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-6' to each side of the equation. 6 + -6 + 2x = 80 + -6 Combine like terms: 6 + -6 = 0 0 + 2x = 80 + -6 2x = 80 + -6 Combine like terms: 80 + -6 = 74 2x = 74 Divide each side by '2'. x = 37 Simplifying x = 37

OpenStudy (anonymous):

Yes??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

em

OpenStudy (anonymous):

Did you do that? Or was it a solver or something else?

OpenStudy (anonymous):

That does not look like the work of a human

OpenStudy (anonymous):

Much wow. You blocked me for taking time out of my day to try to help you. You are trying to do problems that you cannot even begin to do, well out of your math abilities, and you use a solver to do your work for you

OpenStudy (anonymous):

BYE I NEVER TOLD U TO HELP ME

OpenStudy (anonymous):

Its non of your business ok you dont even help at all ok there other that make it better to understand then u you just a wast of time and yeah duh ik I dont know the answer that why I ask for help ****** s

OpenStudy (anonymous):

Right mf